OpenStudy (anonymous):

Verify: (cot x)/(1 + csc x) = (csc x - 1)/(cot x) I realized that if I cross-multiplied: (cot^2 x) = (csc^2 x - 1) (cot^2 x + 1) = csc^2 x Am I able to do this? Or is this technically incorrect?

OpenStudy (anonymous):

you will be able to do

OpenStudy (anonymous):

multiply the numerator and denominator by cosec x -1 and then simplify

OpenStudy (anonymous):

cot x/ (cosex +1) = cotx *( (cosex -1) / ( (cosex -1)* (cosex +1))

OpenStudy (anonymous):

= cotx *(cosecx -1) /(cosec^x-1 ) =cotx *(cosecx -1) / (cotx)^2 =(csc x - 1)/(cot x)

OpenStudy (anonymous):

Okay. Thank you!

OpenStudy (anonymous):

yw